OpenStudy (anonymous):

A square with sides of 3sqrt2 is inscribed in a circle. What is the area of one of the sectors formed by the radii to the vertices of the square? help meeee

OpenStudy (anonymous):

A side of the square is 3*sqrt(2) So, using pythagoras, we can find the diagonal of the square (since the diagonal forms a right angle with the square's sides): C^2=A^2+B^2 - in our case A and B are 3*sqrt(2) ==> C^2 = 18 + 18 = 36 ==> C = 6 The diagonal of the square is 6, and that is also the diameter of the circle. The radius of the circle will be half the diameter. ==> r=3 The area of the entire circle will be pi * r^2 = 9 * pi So, one sector of the circle (one slice) will be a fourth of the entire area: The answer will be: (9/4)*pi

OpenStudy (anonymous):

For such an inscribed square, the radius "R" of the circle is given by the formula a² R² = ----- where a is the length of the side of the rectangle 2 In this case (3√2)² 9*2 R² = ------ = ----- = 9 2 2 so R = √9 => R = 3

OpenStudy (anonymous):

Now radii to the vertices form an angle of 90° so we have to find the area of a sector of angle 90° angle 90° 22 area of sector = -------- * πR² = ------ * ----- * 3² 360° 360° 7 1 22 22 * 9 198 = ----- * ----- * 9 = -------- = ------- = 7.07 sq units 4 7 4 * 7 28